The dipol moment of LIh is 1.9×10^-29cm. and interatomic distance between li and h is 1.56 A. What is %ionic character of lih

Grade:12

1 Answers

Arun
6 years ago

We have, μmolecule = δ × d

Or,                          1.9 × 10 –29 = δ × 1.56 × 10 –10

Therefore,              δ = 1.23 × 10 –19 C

Therefore,          % of ionic nature = (1.23 × 10 –19)/(1.602 × 10–19) × 100 %

= 76.82%